The molecular formula of a compound used for bleaching hair is h2o2. Calculate (a) mass of this compound that would contain 2.5 moles (b) no.of moles of this compound that would exactly weigh 30g .
1
Expert's answer
2016-04-29T09:24:03-0400
a) m(H2O2) = n(H2O2) ∙ M(H2O2) = 2.5 mol ∙ 34 g/mol = 85 g
b) n(H2O2) = m(H2O2) / M(H2O2) = 30 g / 34 g/mol = 0.88 mol
Answer: a) 85 g, b) 0.88 mol